Faculty weekly salaries for 10 teachers in a certain school were coded by dividing each salary by 100. Find the variance of these salaries if the coded observations are 18, 15, 21, 19, 13, 15, 14, 23, 18 and 16 rupees.


1
Expert's answer
2021-01-26T02:40:16-0500

sample is:

1800, 1500, 2100, 1900, 1300, 1500, 1400, 2300, 1800, 1600

n = 10

The formula for variance:


"\frac{\sum (x_i-\bar x)^2}{n-1}"

"\bar x = \frac{\sum x_i}{n}=\frac{1800+1500+2100+1900+1300+1500+1400+2300+1800+1600}{10}=1720"

"variance=\frac{(1800-1720)^2+(1500-1720)^2+...+(1600-1720)^2}{9}=319.026"


Answer: 319.026